### v2

# Maxitive functions with respect to general orders

Abstract.
In decision-making, maxitive functions are used for worst-case and best-case evaluations. Maxitivity gives rise to a rich structure that is well-studied in the context of the pointwise order. In this article, we investigate maxitivity with respect to general preorders and provide a representation theorem for such functionals. The results are illustrated for different stochastic orders in the literature, including the usual stochastic order, the increasing convex/concave order, and the dispersive order.
Key words: Maxitive functions, representation theorem, stochastic orderings
AMS 2020 Subject Classification: 62C86, 47H05, 28A25, 91B05
The second author was partially supported by Fundación Séneca - ACyT Región de Murcia project 21955/PI/22, Agencia Estatal de Investigación (Government of Spain) and Project PID2021-122126NB-C32 funded by MICIU/AEI /10.13039/ 501100011033/ and FEDER A way of making Europe.
1. Introduction
In the context of decision-making under risk and uncertainty, one usually considers a set of gambles $X$ and a functional $\psi\colon X\to\mathbb{R}$, which assigns to each gamble $f\in X$ a number $\psi(f)$ representing its price, risk, or utility. In the theory of imprecise probabilities, these functionals are called previsions, see e.g. [25], while in financial mathematics they are often referred to as risk measures [9] or nonlinear expectations [18].
In many cases, the functional $\psi$ takes the form of an expectation, where the value $\psi(f)$ represents the weighted average of the values associated with the gamble $f$. However, decisions are often based on assessments of best-case or worst-case scenarios, rather than relying on averages. These evaluations are described with maxitive functions. A functional $\psi$ is called maxitive if $\psi(f\vee g)=\psi(f)\vee\psi(g)$ for any two gambles $f$ and $g$, where $f\vee g$ denotes the pointwise supremum of $f$ and $g$.
Maxitive functionals play a crucial role in possibility theory [7, 8, 28] and serve as aggregation functions in several fields such as economics, information fusion, and control theory [10, 24]. Under reasonable assumptions, maxitive functions can be represented in terms of maxitive integrals, such as the Shilkret integral discussed in [23, 2, 19, 13]. Beyond the theory of decision-making, maxitive functions appear in other contexts, including idempotent or tropical analysis [14], or in large deviations theory [12, 13, 20].
In probability theory and statistics, and particularly in decision theory, there exists a wide range of criteria to define order relations between gambles beyond the pointwise order, see [17, 22]. While the majority of the literature has concentrated on maxitivity with respect to pointwise order, little attention has been paid to maxitivity with respect to general orders. The aim of the present work is to close this gap by studying maxitivity with respect to preorders. This allows, in particular, the application of the maxilive structure and its beneficial properties to larger classes of functions. By passing from the pointwise order to weaker stochastic orders, new maxitive functionals are obtained; e.g., it is shown in [26] that the value at risk is maxitive with respect to the usual stochastic order.
[…]
The paper is organized as follows: Section 2 covers basic concepts and definitions. In Section 3, we present and prove the main result. Section 4 discusses variations of the main result for different stochastic orders. Finally, in an appendix, we provide representation results for maxitive risk measures on $C_{b}$ and auxiliary results for quantile functions.
2. Basic concepts and definitions
Let $X$ be a non-empty set and let $\Phi$ be a non-empty set of functions $\phi\colon X\to\mathbb{R}$. Throughout, the set $X$ is endowed with the preorder
$x\preccurlyeq y\quad\mbox{if and only if}\quad\phi(x)\leq\phi(y)\quad\mbox{forall }\phi\in\Phi.$ (2.1)
Further, let $\mathcal{Y}$ be a family of subsets of $X$ which includes all finite subsets. Typical examples of $\mathcal{Y}$ are the family of all finite subsets, the family of all at most countable subsets, or the power set.
Definition 2.1.
$(X,\preccurlyeq)$ is $\mathcal{Y}$-complete if every $Y\in\mathcal{Y}$ which is bounded from above has a unique supremum $\vee Y$.
Definition 2.2.
Let $\psi\colon X\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$ given by $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$.
  • (1)
    The function $\psi$ is monotone if $x\preccurlyeq y$ implies $\psi(x)\leq\psi(y)$.
  • (2)
    If $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete, then the function $\psi$ is $\mathcal{Y}$-maxitive if for every $Y\in\mathcal{Y}$,
    $\sup_{y\in Y}\psi(y)=\begin{cases}\psi(\vee Y),&\mbox{ if }Y\mbox{ is bounded from above,}\\ \infty,&\mbox{ else.}\end{cases}$ (2.2)
  • (3)
    The function $\psi$ has (upwards) $\mathcal{Y}$-directed sublevel sets if for each $s\in\mathbb{R}$, the set $A_{s}$ is $\mathcal{Y}$-directed, i.e., for every $Y\in\mathcal{Y}$ with $Y\subset A_{s}$, there exists $y^{\ast}\in A_{s}$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$.
These concepts are related as follows.
Lemma 2.3.
Suppose that $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete. Then a function $\psi\colon X\to(-\infty,\infty]$ is $\mathcal{Y}$-maxitive if and only if $\psi$ is monotone and has $\mathcal{Y}$-directed sublevel sets.
Proof.
First, suppose that $\psi$ is $\mathcal{Y}$-maxitive. Then for $x,y\in X$ with $x\preccurlyeq y$, it follows from $\mathcal{Y}$-maxitivity with $Y:=\{x,y\}$ that $\psi(x)\leq\psi(x\vee y)=\psi(y)$. Moreover, for $Y\in\mathcal{Y}$ with $Y\subset A_{s}$, it holds $\sup_{y\in Y}\psi(y)\leq s$ and $\mathcal{Y}$-maxitivity implies that $Y$ is bounded from above with $\vee Y\in A_{s}$.
Second, if $\psi$ is monotone, then the right hand side of (2.2) is greater than the left hand side. If $\psi$ has $\mathcal{Y}$-directed sublevel sets, then either $\sup_{y\in Y}\psi(y)=\infty$ or $\sup_{y\in Y}\psi(y)=s<\infty$ in which case $\vee Y\in A_{s}$ so that $\psi(\vee Y)\leq s$, i.e., the right hand side of (2.2) is smaller than the left hand side. ∎
As a typical example, we consider the set $X=[-\infty,\infty]^{\Omega}$ of all [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued functions on a non-empty set $\Omega$, and the family $\Phi:=\{\phi_{\omega}\colon\omega\in\Omega\}$ where $\phi_{\omega}\colon X\to\mathbb{R}$ is given by $\phi_{\omega}(f)=\arctan(f(\omega))$. [^1] We set $\arctan(\infty):=\pi/2$ and $\arctan(-\infty):=-\pi/2$. Then $\Phi$ induces the pointwise order on $X$, and the concept of $\mathcal{Y}$-maxitivity amounts to the usual concepts of maxitivity studied in the literature, see e.g. [2] and references therein. In particular, if $\mathcal{Y}$ is the family of all finite subsets of $X$, then the set of all $\mathcal{Y}$-maxitive functions is exactly the set of all maxitive functions. Likewise, if $\mathcal{Y}$ is the family of countable subsets of $X$ (respectively the family of all subsets of $X$), then the set of all $\mathcal{Y}$-maxitive functions is the set of all countably maxitive functions (respectively the set of all completely maxitive functions).
Definition 2.4.
A function $\alpha\colon{\Phi}\to[-\infty,\infty]$ is $\mathcal{Y}$-stable if for every $Y\in\mathcal{Y}$ with $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, there exists $y^{\ast}\in X$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$ and $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$.
Proposition 2.5.
Suppose that $A=\{x\in X\colon\phi(x)\leq\alpha(\phi)\mbox{ for all }\phi\in\Phi\}$ for some $\alpha\colon\Phi\to[-\infty,\infty]$. Then
$A=\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\},$
where $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ for all $\phi\in\Phi$. In this case, $A$ is $\mathcal{Y}$-directed, if and only if $\alpha$ is $\mathcal{Y}$-stable, if and only if $\alpha_{\min}$ is $\mathcal{Y}$-stable.
Proof.
First, since $\phi(x)\leq\alpha(\phi)$ for all $x\in A$ and $\phi\in\Phi$, it holds $\alpha_{\min}(\phi)\leq\alpha(\phi)$ for all $\phi\in\Phi$, which shows that $\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}\subset A$. Conversely, if $x\in A$, it holds $\phi(x)\leq\alpha_{\min}(\phi)$ for all $\phi\in\Phi$, which shows that $A\subset\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}$.
Second, if $A$ is $\mathcal{Y}$-directed and $Y\in\mathcal{Y}$ satisfies $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, then $Y\subset A$ so that there exists $y^{\ast}\in A$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$, and therefore $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$, i.e., $\alpha$ is $\mathcal{Y}$-stable. Conversely, if $\alpha$ is $\mathcal{Y}$-stable and $Y\in\mathcal{Y}$ satisfies $Y\subset A$, then $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$ and $\mathcal{Y}$-stability implies that $Y$ is bounded from above by some $y^{\ast}\in A$, i.e., $A$ is $\mathcal{Y}$-directed. The argumentation for $\alpha_{\min}$ is analogous. ∎
Remark 2.6.
In general, for an arbitrary $\mathcal{Y}$-directed subset $A\subset X$, it does not follow that $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ is $\mathcal{Y}$-stable, even if $X$ is $\mathcal{Y}$-complete.
To give a counter example, let $\mathcal{Y}$ be the family of all at most countable subsets, and consider the set $X$ of all functions $f\colon\mathbb{R}\to[0,1]$ such that there exists a bounded interval $I\subset\mathbb{R}$ for which $\{x\in\mathbb{R}\setminus I\colon f(x)>0\}$ is at most countable. We consider the family $\Phi:=\{\phi_{x}\colon x\in\mathbb{R}\}$, where $\phi_{x}(f)=f(x)$. The ordered set $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete. Now, let $A$ be the set of all functions $f\in X$ such that $\{x\in\mathbb{R}\colon f(x)>0\}$ is at most countable. Then the set $A$ is $\mathcal{Y}$-directed and $\alpha_{\min}(\phi_{x})=\sup_{f\in A}f(x)=1$ for all $x\in\mathbb{R}$. To show that $\alpha_{\min}$ is not $\mathcal{Y}$-stable, consider for each $n\in\mathbb{N}$ the indicator function $1_{[-n,n]}$ which assumes the value $1$ on $[-n,n]$ and 00 on $[-n,n]^{c}$. Then it holds $\sup_{n\in\mathbb{N}}1_{[-n,n]}(x)\leq\alpha_{\min}(\phi_{x})$ for all $x\in\mathbb{R}$, but there is no function $f\in X$ with $1_{[-n,n]}(x)\leq f(x)$ for all $x\in\mathbb{R}$.
3. Main result
As before, let $X$ be a non-empty set, and let $\Phi$ be a non-empty set of functions $\phi\colon X\to\mathbb{R}$ which induces the preorder $x\preccurlyeq y$ if and only if $\phi(x)\leq\phi(y)$ for all $\phi\in\Phi$. From now on, if not stated otherwise, we assume in addition that
  • the set $\Phi$ is endowed with a topology such that $\Phi$ is a Polish space, and the functions $f_{x}\colon\Phi\to\mathbb{R}$ defined by $f_{x}(\phi):=\phi(x)$ are lower semicontinuous for all $x\in X$;
  • $\mathcal{Y}$ is the family of all at most countable subsets.
In that case, we denote the properties $\mathcal{Y}$-complete, $\mathcal{Y}$-maxitive, $\mathcal{Y}$-directed and $\mathcal{Y}$-stable as countably complete, countably maxitive, countably directed and countably stable, respectively. Note that we do not assume that $(X,\preccurlyeq)$ is countably complete.
The following is our main result.
Theorem 3.1.
Let $\psi\colon X\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$ given by $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$. Then the following statements are equivalent:
  • (i)
    $\psi$ is monotone and has countably directed sublevel sets.
  • (ii)
    There exists a function $\alpha\colon\mathbb{R}\times\Phi\to[-\infty,\infty]$ which is non-decreasing in the first argument and countably stable in the second argument such that
    $\psi(x)=\inf\left\{s\in\mathbb{R}\colon\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha(s,\phi)\}\leq 0\right\}\quad\mbox{for all }x\in X.$ (3.1)
    The infimum is attained for all $x\in X$ with $\psi(x)<\infty$.
  • (iii)
    The function $\alpha_{\min}\colon\mathbb{R}\times\Phi\to[-\infty,\infty]$, $\alpha_{\min}(s,\phi):=\sup_{x\in A_{s}}\phi(x)$ is countably stable in the second argument and
    $\psi(x)=\inf\left\{s\in\mathbb{R}\colon\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0\right\}\quad\mbox{for all }x\in X.$ (3.2)
    The infimum is attained for all $x\in X$ with $\psi(x)<\infty$.
  • (iv)
    The function $\alpha^{+}_{\min}\colon\mathbb{R}\times\Phi\to[-\infty,\infty]$, $\alpha_{\min}^{+}(s,\phi):=\inf_{s<t}\alpha_{\min}(t,\phi)$ is countably stable in the second argument and
    $\psi(x)=\inf\left\{s\in\mathbb{R}\colon\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha^{+}_{\min}(s,\phi)\}\leq 0\right\}\quad\mbox{for all }x\in X.$ (3.3)
    The infimum is attained for all $x\in X$ with $\psi(x)<\infty$.
  • (v)
    The function $\alpha_{\min}^{+}$ is countably stable in the second argument and
    $\psi(x)=\sup_{\phi\in\Phi}G(\phi(x),\phi)\quad\mbox{for all }x\in X,$ (3.4)
    where $G\colon\mathbb{R}\times\Phi\to[-\infty,\infty]$, $G(t,\phi):=\inf\{s\in\mathbb{R}\colon t\leq\alpha^{+}_{\min}(s,\phi)\}$.
Proof.
Let $L(\Phi)$ denote the set of all lower semicontinuous functions $f\colon\Phi\to\mathbb{R}$ endowed with the pointwise order. Recall that $f_{x}\colon\Phi\to\mathbb{R}$ is defined by $f_{x}(\phi):=\phi(x)$ for all $x\in X$. Moreover, let $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$ for all $s\in\mathbb{R}$.
[…]
$F(s,f):=\inf_{x\in A_{s}}\sup_{\phi\in\Phi}\{f(\phi)-\phi(x)\}\quad\mbox{for all }s\in I\mbox{ and }f\in L(\Phi).$ (3.5)
Then the mapping $F\colon I\times L(\Phi)\to[-\infty,\infty]$ is well-defined and satisfies the following properties. For every $f,g\in L(\Phi)$ and $s,t\in I$, it holds
  • (1)
    $F(s,f)\leq F(t,f)$ for all $t\leq s$,
  • (2)
    $F(s,f)\leq F(s,g)$ whenever $f\leq g$,
  • (3)
    $F(s,f+c)=F(s,f)+c$ for all $c\in\mathbb{R}$,
  • (4)
    $F(s,\sup_{n\in\mathbb{N}}f_{n})=\sup_{n\in\mathbb{N}}F(s,f_{n})$ for every sequence $(f_{n})_{n\in\mathbb{N}}\subset L(\Phi)$ with $\sup_{n\in\mathbb{N}}f_{n}\in L(\Phi)$,
  • (5)
    $F(s,f)\in(-\infty,\infty]$ and $F(s,0)\in\mathbb{R}$,
  • (6)
    $A_{s}=\{x\in X\colon F(s,f_{x})\leq 0\}$,
  • (7)
    $\psi(x)=\inf\{s\in I\colon F(s,f_{x})\leq 0\}$ for all $x\in X$, and the infimum is attained whenever $\psi(x)<\infty$.
The properties (1)–(3) follow directly from the definition of $F$.
To show (4), let $(f_{n})_{n\in\mathbb{N}}\subset L(\Phi)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in L(\Phi)$. Then it follows from condition (2) that $\sup_{n\in\mathbb{N}}F(s,f_{n})\leq F(s,f)$. By contradiction, we assume there exists $c\in\mathbb{R}$ with
$\sup_{n\in\mathbb{N}}F(s,f_{n})<c<F(s,f).$
Then for every $n\in\mathbb{N}$ there exists $x_{n}\in A_{s}$ with $f_{n}(\phi)-\phi(x_{n})<c$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$. Using (2.1), we obtain $f_{n}(\phi)-\phi(x^{\ast})<c$ for all $\phi\in\Phi$ and $n\in\mathbb{N}$. Taking the supremum first over all $n\in\mathbb{N}$, and second over all $\phi\in\Phi$, we obtain
$\sup_{\phi\in\Phi}\{f(\phi)-\phi(x^{\ast})\}\leq c<F(s,f)=\inf_{x\in A_{s}}\sup_{\phi\in\Phi}\{f(\phi)-\phi(x)\},$
which is a contradiction to $x^{\ast}\in A_{s}$.
To verify (5), suppose by contradiction that $F(s,f)=-\infty$. Then for $\phi\in\Phi$ there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $f(\phi)-\phi(x_{n})\leq-n$ for all $n\in\mathbb{N}$. Since $A_{s}$ is countably directed, there is $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore
$n+f(\phi)\leq\phi(x^{\ast})\quad\mbox{ for all }n\in\mathbb{N}.$
Letting $n\to\infty$ yields $\phi(x^{\ast})=\infty$ in contrast to the assumption that $\phi$ is real-valued. This shows that $F(s,f)\in(-\infty,\infty]$. In particular, we obtain for every $x\in A_{s}$,
$-\infty<F(s,0)=\inf_{y\in A_{s}}\sup_{\phi\in\Phi}\{-\phi(y)\}\leq-\inf_{\phi\in\Phi}\phi(x)<\infty,$
where the last inequality follows from the boundedness assumption on the family $\Phi$, i.e., $F(s,0)\in\mathbb{R}$.
Next, we show (6). If $x\in A_{s}$, then
$F(s,f_{x})$ $=\inf_{y\in A_{s}}\sup_{\phi\in\Phi}\{\phi(x)-\phi(y)\}\leq\sup_{\phi\in\Phi}\{\phi(x)-\phi(x)\}=0.$
If $x\in X$ satisfies $F(s,f_{x})\leq 0$, then there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $\phi(x)-\phi(x_{n})<1/n$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore $\phi(x)-\phi(x^{\ast})\leq 0$ for all $\phi\in\Phi$, which shows $x\preccurlyeq x^{\ast}$. Since $\psi$ is monotone, we conclude that $\psi(x)\leq\psi(x^{\ast})\leq s$, and therefore $x\in A_{s}$.
To show (7), we notice that condition (6) implies that for every $x\in X$,
$\psi(x)=\inf\{s\in I\colon x\in A_{s}\}=\inf\{s\in I\colon F(s,f_{x})\leq 0\},$
and the infimum is attained whenever $\psi(x)<\infty$. Hence, we have verified the conditions (1)–(7).
Due to conditions (3)–(5), for each $s\in I$, the function $F(s,\cdot)\colon L(\Phi)\to(-\infty,\infty]$ satisfies the assumptions of Corollary A.3, and we obtain the representation
$F(s,f)=\underset{\phi\in\Phi}{\sup}\{f(\phi)-\alpha(s,\phi)\}\quad\mbox{ for all }s\in I,\>f\in L(\Phi),$ (3.6)
where $\alpha(s,\phi):=\sup_{f\in C_{b}(\Phi)}\{f(\phi)-F(s,f)\}$ for all $s\in I$ and $\phi\in\Phi$. As a result of condition (7), we obtain
$\psi(x)=\inf\Big{\{}s\in\mathbb{R}\colon\sup_{\phi\in\Phi}\{\phi(x)-\alpha(s,\phi)\}\leq 0\Big{\}}\quad\mbox{for all }x\in X,$ (3.7)
where we set $\alpha(s,\phi):=-\infty$ whenever $s\notin I$. Moreover, the infimum is attained whenever $\psi(x)<\infty$. Due to condition (1), the mapping $\alpha$ is non-decreasing in the first argument. For $s\in\mathbb{R}$, the representation (3.7) implies that $x\in A_{s}$ if and only if $\phi(x)\leq\alpha(s,\phi)$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, it follows from Proposition 2.5 that $\alpha$ is countably stable in the second argument.
$(ii)\Rightarrow(iii)$: Fix $s\in\mathbb{R}$. The representation (3.1) implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha(s,\phi)\}\leq 0$. Since $\alpha$ is countably stable in the second argument, it follows from Proposition 2.5 that $A_{s}$ is countably directed. Moreover, Proposition 2.5 implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, where $\alpha_{\min}(s,\phi):=\sup_{x\in A_{s}}\phi(x)$, and $\alpha_{\min}$ is countably stable in the second argument.
$(iii)\Rightarrow(iv)$: Fix $s\in\mathbb{R}$. It follows from (iii) together with Proposition 2.5 that $A_{s}$ is countably directed and $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$. Hence, $x\in X$ satisfies $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, if and only if $x\in A_{t}$ for all $t>s$, if and only if $\sup_{t>s}\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(t,\phi)\}\leq 0$, if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha^{+}_{\min}(t,\phi)\}\leq 0$. In particular, it follows from Proposition 2.5 that $\alpha^{+}$ is coutable stable in the second argument.
$(iv)\Rightarrow(v)$: Since $\alpha_{\min}^{+}$ is right-continuous and non-decreasing in the first argument, it holds
$\alpha_{\phi}^{-1}([t,\infty))=\left[G(t,\phi),\infty\right)\quad\mbox{for all}t\in\mathbb{R}\mbox{ and }\phi\in\Phi,$
where $\alpha_{\phi}^{-1}([t,\infty)):=\{s\in\mathbb{R}\colon\alpha_{\min}^{+}(s,\phi)\in[t,\infty)\}$ is the preimage of $[t,\infty)$ under $\alpha_{\min}^{+}$. Fix $x\in X$. Suppose first that $\psi(x)<\infty$. It follows from the representation (3.3) that $s\in[\psi(x),\infty)$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}^{+}(s,\phi)\}\leq 0$, if and only if $\phi(x)\leq\alpha_{\min}^{+}(s,\phi)$ for all $\phi\in\Phi$, if and only if $s\in\alpha_{\phi}^{-1}([\phi(x),\infty))$ for all $\phi\in\Phi$. Hence,
$[\psi(x),\infty)=\bigcap_{\phi\in\Phi}\alpha_{\phi}^{-1}([\phi(x),\infty))=\bigcap_{\phi\in\Phi}\big{[}G(\phi(x),\phi),\infty\big{)}.$
This shows that $\psi(x)=\sup_{\phi\in\Phi}G(\phi(x),\phi)$ for all $x\in X$. In case that $\psi(x)=\infty$, then the representation (3.3) implies that for every $n\in\mathbb{N}$ we can find $\phi_{n}\in\Phi$ such that $n\notin\alpha_{\phi}^{-1}([\phi_{n}(x),\infty))$. Hence, for every $n\in\mathbb{N}$,
$n\notin\bigcap_{\phi\in\Phi}\alpha_{\phi}^{-1}([\phi(x),\infty))=\bigcap_{\phi\in\Phi}\big{[}G(\phi(x),\phi),\infty\big{)}.$
This implies that $\psi(x)=\sup_{\phi\in\Phi}G(\phi(x),\phi)=\infty$.
$(v)\Rightarrow(i)$ Since $\alpha^{+}_{\min}$ is countably stable in the second argument, we argue similar as in the implication from (iii) to (iv) that the level sets of $\psi$ are countably directed. Since $G$ is non-decreasing in the first argument, the functional $\psi$ is monotone. ∎
Remark 3.2.
We point out that the statement of Theorem 3.1 remains valid if we replace countable maxitivity by complete maxitivity and countable stability by complete stability. [^3] I.e., $\mathcal{P}(X)$-maxitivity and $\mathcal{P}(X)$-stability, where $\mathcal{P}(X)$ is the power set. The proof follows with the same arguments and the following modifications. In the implication $(i)\Rightarrow(ii)$, we do not rely on Corollary A.3, but rather extend the definition (3.5) of $F(s,f)$ for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$, where $[-\infty,\infty)^{\Phi}$ denotes the set of all functions from $\Phi$ to [,)[$\infty$,$\infty$)[ - ∞ , ∞ ). Direct verification shows that $F$ satisfies the conditions (1)–(3) on $I\times[-\infty,\infty)^{\Phi}$ and $F(s,\sup_{j\in J}f_{j})=\sup_{j\in J}F(s,f_{j})$ for every family $(f_{j})_{j\in J}$ in $[-\infty,\infty)^{\Phi}$. From the latter property, one obtains the representation (3.6) for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$ for a suitable function $\alpha\colon I\times\Phi\to(-\infty,\infty]$. From there, the argumentation is similar as in the proof of $(i)\Rightarrow(ii)$ in Theorem 3.1.
For the rest of this section, the space $X$ is endowed with an operation $+\colon X\times\mathbb{R}\to X$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the translation property if $\psi(x+c)=\psi(x)+c$ for all $x\in X$ and $c\in\mathbb{R}$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the deviation property if $\psi(x+c)=\psi(x)$ for all $x\in X$ and $c\in\mathbb{R}$.
Corollary 3.3.
Suppose that every $\phi\in\Phi$ has the translation property. Let $\psi\colon X\to(-\infty,\infty]$ be a function with $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$ for all $s\in\mathbb{R}$. Then the following statements are equivalent:
  • (i)
    $\psi$ is monotone and has countably directed sublevel sets.
  • (ii)
    There exists a stable function $\gamma\colon\Phi\to[-\infty,\infty]$ with
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\gamma(\phi)\}\quad\mbox{for all }x\in X.$
  • (iii)
    The mapping $\phi\mapsto\alpha_{\min}(0,\phi):=\sup_{x\in A_{0}}\psi(x)$ is stable and
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(0,\phi)\}\quad\mbox{for all }x\in X.$
Proof.
$(i)\Rightarrow(iii)$: Since $\psi$ has the translation property, it holds $\psi(x)\leq s$ if and only if $\psi(x-s)\leq 0$, so that $A_{s}=A_{0}-s$ for all $s\in\mathbb{R}$. Hence, since every $\phi\in\Phi$ has the translation property, we obtain for every $s\in\mathbb{R}$ and $\phi\in\Phi$,
$\alpha_{\min}(s,\phi)=\sup_{x\in A_{s}}\phi(x)=\sup_{x\in A_{0}}\phi(x+s)=\sup_{x\in A_{0}}\phi(x)+s=\alpha_{\min}(0,\phi)+s.$
It follows from Theorem 3.1 that for every $x\in X$,
$\psi(x)$ $=\inf\left\{s\in\mathbb{R}\colon\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0\right\}$
$=\inf\left\{s\in\mathbb{R}\colon\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha_{\min}(0,\phi)\}\leq s\right\}$
$=\underset{\phi\in\Phi}{\sup}\{\phi(x)-\alpha_{\min}(0,\phi)\}.$
$(iii)\Rightarrow(ii)$ is obvious.
$(ii)\Rightarrow(i)$: That $\psi$ is monotone and has countably directed sublevel sets follows from Theorem 3.1. Moreover, since every $\phi\in\Phi$ has the translation property, for every $x\in X$ and $c\in\mathbb{R}$,
$\psi(x+c)=\sup_{\phi\in\Phi}\{\phi(x+c)-\gamma(\phi)\}=\sup_{\phi\in\Phi}\{\phi(x)+c-\gamma(\phi)\}=\psi(x)+c,$
which shows that $\psi$ has the translation property. ∎
Corollary 3.4.
Suppose that $\psi\colon X\to(-\infty,\infty]$ is monotone and has countably directed sublevel sets. If $\psi$ has the deviation property, then $\psi(x)=\infty$ for all $x\in X$.
Proof.
It follows from Theorem 3.1 that for every $x\in X$,
$\psi(x)=\inf\left\{s\in\mathbb{R}\colon\phi(x)\leq\alpha_{\min}(s,\phi)\mbox{ for all }\phi\in\Phi\right\}.$
By similar arguments as in Corollary 3.3, it holds $\alpha_{\min}(s,\phi)=\alpha_{\min}(0,\phi)$ for all $s\in\mathbb{R}$. Hence, it necessary holds $\psi(x)=\infty$. ∎
4. Examples
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an atomless probability space. We denote by $\mathcal{M}$ the set of all random variables $\xi\colon\Omega\to\mathbb{R}$, where two of them are identified if they have the same law. Further, let $\mathcal{M}_{\infty}$ and $\mathcal{M}_{1}$ be the sets of all random variables in $\mathcal{M}$ which are essentially bounded, and have finite first moments, respectively. In the following, we analyze the main results of the previous section for several stochastic orders. For definitions and further details on stochastic orders we refer to [16, 22].
4.1. Usual stochastic order
For $\xi\in\mathcal{M}$, the quantile function $q_{\xi}\colon(0,1)\to\mathbb{R}$ is defined as the left-inverse of the cumulative distribution function, i.e.,
$q_{\xi}(u):=\inf\{v\in\mathbb{R}\colon u\leq\mathbb{P}(\xi\leq v)\}.$
On the one hand, the quantile function $q_{\xi}$ is non-decreasing and left-continuous, see Lemma B.1. On the other hand, every function $q\colon(0,1)\to\mathbb{R}$ which is non-decreasing and left-continuous satisfies $q=q_{\xi}$ for some $\xi\in\mathcal{M}$, see Lemma B.3. Hence, there is a one-to-one relation between the set $\mathcal{M}$ and quantile functions.
The usual stochastic order on $\mathcal{M}$ is defined by
$\xi\preccurlyeq_{\rm st}\eta\quad\mbox{if and only if}\quad q_{\xi}(u)\leq q_{\eta}(u)\mbox{ for all }u\in(0,1).$
The ordered space $(\mathcal{M},\preccurlyeq_{\rm st})$ is a complete lattice, see [11, 17]. In particular, every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ which is bounded from above has a supremum $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ in $\mathcal{M}$ which is determined by $q_{\xi^{\ast}}(u)=\sup_{n\in\mathbb{N}}q_{\xi_{n}}(u)$ for all $u\in(0,1)$.
Let $\Phi_{\rm st}:=\{\phi_{u}\colon u\in(0,1)\}$ denote the family of functions $\phi_{u}\colon\mathcal{M}\to\mathbb{R}$ given by $\phi_{u}(\xi):=q_{\xi}(u)$. We identify $\Phi_{\rm st}$ with $(0,1)$, and endow $(0,1)$ with the Euclidean topology.444$(0,1)$ is not complete with the Euclidean distance. However, it is completely metrizable, that is, there exists an equivalent metric on $(0,1)$ such that $(0,1)$ with the relative Euclidean topology is a complete metric space space. In fact, every countable intersection of open subsets of a complete metric space is completely metrizable, see [27, Theorem 24.12]. Then it holds $\xi\preccurlyeq_{\rm st}\eta$ if and only if $\phi_{u}(\xi)\leq\phi_{u}(\eta)$ for all $u\in(0,1)$. Moreover, for every $\xi\in\mathcal{M}$, the mapping $u\mapsto\phi_{u}(\xi)$ is lower semicontinous. Therefore, the partially ordered space $(\mathcal{M},\preccurlyeq_{\rm st})$ satisfies the assumptions of Section 3.
Lemma 4.1.
Every function $\gamma\colon(0,1)\to[-\infty,\infty)$ is countably stable.
Proof.
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ with $q^{\ast}(u):=\sup_{n\in\mathbb{N}}q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Then the mapping $u\mapsto q^{\ast}(u)$ is real-valued, non-decreasing and left-continuous. Hence, there exists $\xi^{\ast}\in\mathcal{M}$ with $q^{\ast}(u)=q_{\xi^{\ast}}(u)$ for all $u\in(0,1)$. In particular, $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ and $\phi_{u}(\xi^{\ast})=q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, so that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the usual stochastic order.
Theorem 4.2.
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
  • (ii)
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm st}$.
  • (iii)
    There exists a function $\alpha\colon\mathbb{R}\times(0,1)\to[-\infty,\infty)$ which is non-decreasing in the first argument such that
    $\psi(\xi)=\inf\left\{s\in\mathbb{R}\colon\sup_{u\in(0,1)}\{q_{\xi}(u)-\alpha(s,u)\}\leq 0\right\}\quad\mbox{for all }\xi\in\mathcal{M}.$
    The infimum is attained for all $\xi\in\mathcal{M}$ with $\psi(\xi)<\infty$.
  • (iv)
    The mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}G(q_{\xi}(u),u)\quad\mbox{ for all }\xi\in\mathcal{M},$
    where $G(t,u):=\inf\{s\in\mathbb{R}\colon t\leq\alpha_{\min}^{+}(s,u)\}$ for all $(t,u)\in\mathbb{R}\times(0,1)$.
If $\psi$ also has the translation property, then we get the following version of the previous result. We notice that $q_{\xi+c}=q_{\xi}+c$ for all $\xi\in\mathcal{M}$ and $c\in\mathbb{R}$, i.e., the mapping $\phi_{u}\colon\mathcal{M}\to\mathbb{R}$ has the translation property for all $u\in(0,1)$. Hence, we can apply Corollary 3.3 and obtain the following result.
Corollary 4.3.
Let $\psi\colon\mathcal{M}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm st}$.
  • (ii)
    There exists a function $\alpha\colon\mathbb{R}\times(0,1)\to[-\infty,\infty)$ such that
    $\psi(\xi)=\sup_{u\in(0,1)}\{q_{\xi}(u)-\alpha(u)\}\quad\mbox{for all }\xi\in\mathcal{M}.$
  • (iii)
    The mapping $u\mapsto\alpha_{\min}(0,u)=\sup_{\xi\in A_{0}}q_{\xi}(u)$ is real-valued, and
    $\psi(\xi)=\sup_{u\in(0,1)}\{q_{\xi}(u)-\alpha_{\min}(0,u)\}\quad\mbox{ for all}\xi\in\mathcal{M}.$
Recall that the value at risk ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ at level $u\in(0,1)$, which is defined by ${\rm VaR}_{u}(\xi):=q_{\xi}(u)$, is a monetary risk measure. A detailed discussion on risk measures can be found in [9]. The previous Corollary 4.3 is closely related to [26, Theorem 4] which provides a representation result for lower semicontinuous complete maxitive functionals $\mathcal{M}_{1}\to\mathbb{R}$ having the translation property. A related representation result was also shown in [1, Proposition 4] for the class of benchmark-loss VaR which were introduced by Bignozzi et al. [1].
Remark 4.4.
For every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$ such that $\vee_{n\in\mathbb{N}}\xi_{n}$ exists, it holds ${\rm VaR}_{u}(\vee_{n\in\mathbb{N}}\xi_{n})=\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})$. Nevertheless, the function ${\rm VaR}_{u}\colon\mathcal{M}\to\mathbb{R}$ is not countably maxitive because it may happen that $\sup_{n\in\mathbb{N}}{\rm VaR}_{u}(\xi_{n})<\infty$ for an unbounded sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}$. However, the function ${\rm VaR}_{u}$ can also be defined on the set $\mathcal{M}^{\ast}$ of all extended random variables $\xi\colon\Omega\to(-\infty,\infty]$. Then the supremum $\vee_{i\in I}\xi_{i}$ exists for every family $(\xi_{i})_{i\in I}\subset\mathcal{M}^{\ast}$ and ${\rm VaR}_{u}\colon\mathcal{M}^{\ast}\to(-\infty,\infty]$ is completely maxitive.
4.2. Increasing convex stochastic order
For $\xi\in\mathcal{M}_{1}$, we define the integrated quantile function $Q_{\xi}\colon(0,1)\to\mathbb{R}$ by
$Q_{\xi}(u):=\int_{u}^{1}q_{\xi}(v)\,{\rm d}v.$
The function $u\mapsto Q_{\xi}(u)$ is concave, with $Q_{\xi}(0^{+})=\mathbb{E}[\xi]$ and $Q_{\xi}(1^{-})=0$. Conversely, every concave function $Q\colon(0,1)\to\mathbb{R}$ with $Q(0^{+})\in\mathbb{R}$ and $Q(1^{-})=0$ is of the form $Q=Q_{\xi}$ for some random variable $\xi\in\mathcal{M}_{1}$, see Lemma B.4. The increasing convex order on $\mathcal{M}_{1}$ is defined by
$\xi\preccurlyeq_{\rm icx}\eta\quad\mbox{if and only if}\quad Q_{\xi}(u)\leq Q_{\eta}(u)\quad\mbox{ for all }u\in(0,1).$
The ordered space $(\mathcal{M}_{1},\preccurlyeq_{\rm icx})$ is a complete lattice, see [11, 17]. In particular, every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{1}$ which is bounded from above has a supremum $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ which is determined by
$Q_{\xi^{\ast}}(u)={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)(u)\quad\mbox{ forall }u\in(0,1).$
Here, ${\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)$ denotes the concave hull of the sequence $(Q_{\xi_{n}})_{n\in\mathbb{N}}$, i.e., the smallest concave function $Q^{\ast}\colon(0,1)\to[-\infty,\infty]$ with $\sup_{n\in\mathbb{N}}Q_{\xi_{n}}(u)\leq Q^{\ast}(u)$ for all $u\in(0,1)$. [^5] Notice that ${\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)$ is the pointwise infimum over all concave functions $Q\colon(0,1)\to(-\infty,\infty]$ with $Q_{\xi_{n}}(u)\leq Q(u)$ for all $u\in(0,1)$ and $n\in\mathbb{N}$.
Let $\Phi_{\rm icx}:=\{\phi_{u}\colon u\in(0,1)\}$ be the family of functions $\phi_{u}\colon\mathcal{M}_{1}\to\mathbb{R}$ given by $\phi_{u}(\xi):=Q_{\xi}(u)$. We identify $\Phi_{\rm icx}$ with $(0,1)$, and endow $(0,1)$ with the Euclidean topology. For every $\xi\in\mathcal{M}_{1}$, the function $u\mapsto\phi_{u}(\xi)$ is continuous. Hence, the partially ordered space $(\mathcal{M}_{1},\preccurlyeq_{\rm icx})$ satisfies the assumptions of Section 3.
Lemma 4.5.
Every concave function $\gamma\colon(0,1)\to[-\infty,\infty)$ is countably stable.
Proof.
Let $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{1}$ with $\vee_{n\in\mathbb{N}}Q_{\xi_{n}}(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $\gamma$ is concave, it holds $Q^{\ast}(u):={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}},\cdots)(u)\leq\gamma(u)$ for all $u\in(0,1)$. Since $u\mapsto Q^{\ast}(u)$ is real-valued, we obtain $Q_{\xi^{\ast}}=Q^{\ast}$ where $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}\in\mathcal{M}_{1}$, and $Q_{\xi^{\ast}}(u)=Q^{\ast}(u)\leq\gamma(u)$ for all $u\in(0,1)$. ∎
For a function $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, from which it is straightforward to verify that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued and concave for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. [^6] As for the concavity, we notice that for $\xi_{1},\xi_{2}\in A_{s}$, there exists $\xi^{\ast}\in A_{s}$ with $Q_{\xi^{\ast}}={\rm conc}(Q_{\xi_{1}},Q_{\xi_{2}})$ such that $\lambda Q_{\xi_{1}}(u_{1})+(1-\lambda)Q_{\xi_{2}}(u_{2})\leq Q_{\xi^{\ast}}(\lambda u_{1}+(1-\lambda)u_{2})$ for all $u_{1},u_{2},\lambda\in(0,1)$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the increasing convex stochastic order.
Theorem 4.6.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
  • (ii)
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm icx}$.
  • (iii)
    There exists a function $\alpha\colon\mathbb{R}\times(0,1)\to[-\infty,\infty)$ which is non-decreasing in the first argument and concave in the second argument such that
    $\psi(\xi)=\inf\left\{s\in\mathbb{R}\colon\sup_{u\in(0,1)}\{Q_{\xi}(u)-\alpha(s,u)\}\leq 0\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
    The infimum is attained for all $\xi\in\mathcal{M}_{1}$ with $\psi(\xi)<\infty$.
  • (iv)
    The mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued and concave for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}G(Q_{\xi}(u),u)\quad\mbox{ for all }\xi\in\mathcal{M}_{1},$
    where $G(t,u):=\inf\{s\in\mathbb{R}\colon t\leq\alpha_{\min}^{+}(s,u)\}$ for all $(t,u)\in\mathbb{R}\times(0,1)$.
If $\psi$ in addition has the translation property, then we can argue as follows. Since the function $\xi\mapsto Q_{\xi}$ does not have the translation property, we consider the modified family $\Phi^{\prime}_{\rm icx}:=\{{\rm ES}_{u}\colon u\in(0,1)\}$. Here, ${\rm ES}_{u}\colon\mathcal{M}_{1}\to\mathbb{R}$ is the expected shortfall given by
${\rm ES}_{u}(\xi):=\frac{1}{1-u}Q_{\xi}(u)\quad\mbox{ for all }u\in(0,1).$
The expected shortfall is a convex risk measure, see e.g. [9] for further details. In particular, the functions ${\rm ES}_{u}$ have the translation property and also induce the partial order $\preccurlyeq_{\rm icx}$. However, by changing to the family $\Phi^{\prime}_{\rm icx}$, we also change the set of countably stable functions. More precisely, a function $\gamma\colon(0,1)\to[-\infty,\infty]$ is countably stable w.r.t. $\Phi_{\rm icx}$ if and only if the function $u\mapsto\frac{1}{1-u}\gamma(u)$ is countably stable w.r.t. $\Phi_{\rm icx}^{\prime}$. Moreover, the minimal penalty function $\alpha_{\min}^{\prime}$ w.r.t. $\Phi_{\rm icx}^{\prime}$ is given by
$\alpha_{\min}^{\prime}(s,u)=\tfrac{1}{1-u}\alpha_{\min}(s,u).$
As a consequence of Corollary 3.3, we obtain the following result.
Corollary 4.7.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
  • (ii)
    There exists a concave function $\alpha\colon(0,1)\to[-\infty,\infty)$ such that
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{{\rm ES}_{u}(\xi)-\tfrac{1}{1-u}\alpha(u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
  • (iii)
    The mapping $u\mapsto\alpha_{\min}(0,u)$ is real-valued and concave. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{{\rm ES}_{u}(\xi)-\tfrac{1}{1-u}\alpha_{\min}(0,u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
The previous Corollary 4.7 is closely related to [26, Theorem 5]. Moreover, it is shown in [26] that the function ${\rm ES}_{u}$ does not always preserve the maximum of two randoms variables w.r.t. $\preccurlyeq_{\rm icx}$, i.e., the function ${\rm ES}_{u}$ is not maxitive w.r.t. $\preccurlyeq_{\rm icx}$.
4.3. Convex stochastic order
The convex order on $\mathcal{M}_{1}$ is defined by
$\xi\preccurlyeq_{\rm cx}\eta\quad\mbox{if and only if}\quad\mathbb{E}[\xi]=\mathbb{E}[\eta]\mbox{ and }\xi\preccurlyeq_{\rm icx}\eta.$
The partially ordered space $(\mathcal{M}_{1},\preccurlyeq_{\rm cx})$ is not a lattice on $\mathcal{M}_{1}$ because it is not possible to define the maximum of two random variables $\xi,\nu\in\mathcal{M}_{1}$ for which $\mathbb{E}[\xi]\neq\mathbb{E}[\eta]$. However, for $r\in\mathbb{R}$, the convex order agrees with the increasing convex order on the set
${\mathcal{M}}_{1}^{r}:=\left\{\xi\in{\mathcal{M}}_{1}\colon\mathbb{E}[\xi]=r\right\}.$
In particular, $(\mathcal{M}^{r}_{1},\preccurlyeq_{\rm cx})$ is a complete lattice. For $\psi\colon\mathcal{M}^{r}_{1}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}^{r}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}\colon\mathbb{R}\times(0,1)\to[-\infty,\infty]$ by $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}Q_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the convex stochastic order.
Theorem 4.8.
Let $\psi\colon\mathcal{M}_{1}^{r}\to\mathbb{R}$ be a function. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm cx}$.
  • (ii)
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm cx}$.
  • (iii)
    There exists a function $\alpha\colon\mathbb{R}\times(0,1)\to[-\infty,\infty)$ which is non-decreasing in the first argument and concave in the second argument such that
    $\psi(\xi)=\min\left\{s\in\mathbb{R}\colon\sup_{u\in(0,1)}\{Q_{\xi}(u)-\alpha(s,u)\}\leq 0\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}^{r}.$
  • (iv)
    The mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued and concave for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}G(Q_{\xi}(u),u)\quad\mbox{ for all }\xi\in\mathcal{M}_{1}^{r},$
    where $G(t,u):=\inf\{s\in\mathbb{R}\colon t\leq\alpha_{\min}^{+}(s,u)\}$ for all $(t,u)\in\mathbb{R}\times(0,1)$.
4.4. Increasing concave stochastic order
For $\xi\in\mathcal{M}_{1}$, we define the function $\bar{Q}_{\xi}\colon(0,1)\to\mathbb{R}$ by
$\bar{Q}_{\xi}(u):=-{Q}_{-\xi}(u)=-\int_{u}^{1}q_{-\xi}(v)\,{\rm d}v.$
The mapping $t\mapsto\bar{Q}_{\xi}(t)$ is convex and satisfies $\bar{Q}_{\xi}(0^{+})=\mathbb{E}[\xi]$ and $\bar{Q}_{\xi}(1^{-})=0$. Conversely, every convex function $\bar{Q}\colon(0,1)\to\mathbb{R}$ with $\bar{Q}(0^{+})\in\mathbb{R}$ and $\bar{Q}(1^{-})=0$ is of the form $\bar{Q}=\bar{Q}_{\xi}$ for some random variable $\xi\in\mathcal{M}_{1}$. The increasing concave stochastic order is defined by
$\xi\preccurlyeq_{\rm icv}\eta\quad\mbox{if and only if }\quad-\eta\preccurlyeq_{\rm icx}-\xi.$
That is, $\xi\preccurlyeq_{\rm icv}\eta$ if and only if $\overline{Q}_{\xi}(u)\leq\overline{Q}_{\eta}(u)$ for all $u\in(0,1)$. The partially ordered space $(\mathcal{M}_{1},\preccurlyeq_{\rm icv})$ is a complete lattice, see [11, 17]. In particular, every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{1}$ which is bounded from above w.r.t. $\preccurlyeq_{\rm icv}$ has a supremum $\xi^{\ast}=\vee_{n\in\mathbb{N}}\xi_{n}$ which is determined by $\bar{Q}_{\xi^{\ast}}(u)=\vee_{n\in\mathbb{N}}\bar{Q}_{\xi_{n}}(u)$ for all $u\in(0,1)$. We consider the family $\Phi_{\rm icv}:=\{\phi_{u}\colon u\in(0,1)\}$ of functions $\phi_{u}\colon\mathcal{M}_{1}\to\mathbb{R}$ given by $\phi_{u}(\xi):=\bar{Q}_{\xi}(u)$. Hence, the partially ordered space $(\mathcal{M}_{1},\preccurlyeq_{\rm icv})$ satisfies the conditions of Section 3. With arguments similar to those for the usual stochastic order, we get the following result.
Lemma 4.9.
Every function $\gamma\colon(0,1)\to[-\infty,\infty)$ is countably stable.
For a function $\psi\colon\mathcal{M}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{1}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u):=\sup_{\xi\in A_{s}}\bar{Q}_{\xi}(u)$ and $\alpha^{+}_{\min}(s,u):=\inf_{s<t}\alpha_{\min}(t,u)$ for all $s\in\mathbb{R}$ and $u\in(0,1)$. If $\psi$ is countably maxitive, then the sublevel sets are countably directed, so that the mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. As a direct application of Theorem 3.1 and Lemma 2.3, we obtain the following representation result for maxitive functions w.r.t. the increasing concave stochastic order.
Theorem 4.10.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function. Then the following conditions are equivalent.
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icv}$.
  • (ii)
    $\psi$ is monotone and has countably directed sublevel sets w.r.t. $\preccurlyeq_{\rm icv}$.
  • (ii)
    There exists a function $\alpha\colon\mathbb{R}\times(0,1)\to[-\infty,\infty)$ which is non-decreasing in the first argument such that
    $\psi(\xi)=\inf\left\{s\in\mathbb{R}\colon\sup_{u\in(0,1)}\{\bar{Q}_{\xi}(u)-\alpha(s,u)\}\leq 0\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
    The infimum is attained for all $\xi\in\mathcal{M}_{1}$ with $\psi(\xi)<\infty$.
  • (iii)
    The mapping $u\mapsto\alpha_{\min}(s,u)$ is real-valued for all $s\in\mathbb{R}$ with $A_{s}\neq\emptyset$. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}G\left(\bar{Q}_{\xi}(u),u\right)\quad\mbox{ for all }\xi\in\mathcal{M},$
    where $G(t,u):=\inf\{s\in\mathbb{R}\colon t\leq\alpha_{\min}^{+}(s,u)\}$ for all $(t,u)\in\mathbb{R}\times(0,1)$.
In case that $\psi$ has the translation property, we can proceed as follows. Since the function $\xi\mapsto\bar{Q}_{\xi}$ does not have the translation property, we consider the modified family $\Phi_{icv}^{\prime}:=\{\overline{\rm ES}_{u}\colon u\in(0,1)\}$, where $\overline{{\rm ES}}_{u}\colon\mathcal{M}_{1}\to\mathbb{R}$ is defined by $\overline{{\rm ES}}_{u}(\xi):=-{{\rm ES}}_{u}(-\xi)$. The functions $\overline{{\rm ES}}_{u}$ are translation invariant and induce the order $\preccurlyeq_{\rm icv}$. The families $\Phi_{\rm icv}^{\prime}$ and $\Phi_{\rm icv}$ result in the same countably stable functions. In addition, we obtain
$\alpha_{\min}^{\prime}(s,u)=\frac{1}{1-u}\alpha_{\min}(s,u),$
where $\alpha_{\min}^{\prime}$ is the minimal penalty function w.r.t. $\Phi_{\rm icv}^{\prime}$. As a consequence of Corollary 3.3, we obtain the following result.
Corollary 4.11.
Let $\psi\colon\mathcal{M}_{1}\to(-\infty,\infty]$ be a function with the translation property. Then the following conditions are equivalent:
  • (i)
    $\psi$ is countably maxitive w.r.t. $\preccurlyeq_{\rm icv}$.
  • (ii)
    There exists a function $\alpha\colon(0,1)\to[-\infty,\infty)$ such that
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{\overline{{\rm ES}}_{u}(\xi)-\tfrac{1}{1-u}\alpha(u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
  • (iii)
    The mapping $u\mapsto\alpha_{\min}(0,u)$ is real-valued. Moreover, it holds
    $\psi(\xi)=\sup_{u\in(0,1)}\left\{\overline{{\rm ES}}_{u}(\xi)-\tfrac{1}{1-u}\alpha_{\min}(0,u)\right\}\quad\mbox{for all }\xi\in\mathcal{M}_{1}.$
4.5. Dispersive stochastic order
The dispersive stochastic order on $\mathcal{M}$ is defined by
$\xi\preccurlyeq_{\rm disp}\eta\quad\mbox{if and only if}\quad q_{\xi}(v)-q_{\xi}(u)\leq q_{\eta}(v)-q_{\eta}(u)\mbox{ for all }0<u\leq v<1.$
Remark 4.12.
The dispersive order is not antisymmetric and is therefore not a partial order. However, by considering the quotient space $\mathcal{M}/_{\sim}$ w.r.t. the equivalence relation $\xi{\sim}\eta$ if and only if $\xi=\eta+c$ for some $c\in\mathbb{R}$, we obtain that the partially ordered space $(\mathcal{M}/_{\sim},\preccurlyeq_{\rm disp})$ is a lattice, see [17]. It is shown in Corollary C.4 that the lattice $(\mathcal{M}_{\infty}/_{\sim},\preccurlyeq_{\rm disp})$ is order complete.
It follows from Proposition C.3 that every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{\infty}$ which is bounded from above w.r.t. $\preccurlyeq_{\rm disp}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determined by
$q_{\xi^{\ast}}(u):={\rm TV}_{[0,u]}\left(\{q_{\xi_{n}}\colon n\in\mathbb{N}\}\right)\quad\mbox{ for all }u\in(0,1].$
Here, for $0\leq u<v\leq 1$, we denote by ${\rm TV}_{[u,v]}\left(Q\right):=\sup_{\pi\in\Pi[u,v]}S_{\pi}(Q)$ the total variation along $[u,v]$ of a non-empty set $Q$ of quantile functions, where $\Pi[u,v]$ is the set of all partitions of the form $\pi=\{u=t_{0}<t_{1}<\cdots<t_{n}=v\}$ for some $n\in\mathbb{N}$, for which we set $S_{\pi}(Q):=\sum_{i=1}^{n}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))$ with $q(0):=q(0^{+})$ and $q(1):=q(1^{-})$. It follows from Proposition B.2 that
$\xi\preccurlyeq_{\rm disp}\eta\quad\mbox{if and only if}\quad q_{\xi}(v)-q_{\xi}^{+}(u)\leq q_{\eta}(v)-q_{\eta}^{+}(u)\mbox{ for all }0<u<v<1,$ (4.1)
where $q_{\xi}^{+}(u):=\inf\{x\in\mathbb{R}\colon\mathbb{P}(\xi\leq x)>u\}$ is the right quantile function of $\xi$.
In the following, we consider the family $\Phi_{\rm disp}:=\{\phi_{u,v}\colon 0<u<v<1\}$ of functions $\phi_{u,v}\colon\mathcal{M}_{\infty}\to\mathbb{R}$ defined by $\phi_{u,v}(\xi):=q_{\xi}(v)-q_{\xi}^{+}(u)$. We identify $\Phi_{\rm disp}$ with the set $\Delta:=\{(u,v)\in\mathbb{R}^{2}\colon 0<u<v<1\}$ which we endow with the Euclidean topology. [^7] Although $\Delta$ is not a complete metric space, it is completely metrizable as it is an open subset, see [27, Theorem 24.12]. In particular, $\Delta$ is a Polish space with the Euclidean topology. It follows from (4.1) that the family $\Phi_{\rm disp}$ induces the dispersive order on $\mathcal{M}_{\infty}$. By Proposition B.1, the mapping $(u,v)\mapsto\phi_{u,v}(\xi)$ is lower semicontinuous on $\Delta$ for all $\xi\in\mathcal{M}_{\infty}$. Hence, the ordered space $(\mathcal{M}_{\infty},\preccurlyeq_{\rm disp})$ satisfies the assumptions of Section 3.
Lemma 4.13.
A function $\alpha\colon\Delta\to[0,\infty)$ is countably stable if there exists a non-decreasing function $\beta\colon[0,1]\to\mathbb{R}$ such that $\alpha(u,v)=\beta(v)-\beta(u)$ for all $(u,v)\in\Delta$.
Proof.
Suppose that $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{\infty}$ satisfies $\sup_{n\in\mathbb{N}}\{q_{\xi_{n}}(v)-q_{\xi_{n}}^{+}(u)\}\leq\alpha(u,v)$ for all $(u,v)\in\Delta$. As an application of Proposition C.2 and Proposition C.1, we obtain that
${\rm TV}_{[0,1]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})=\lim_{t\to 1^{-}}{\rm TV}_{[0,t]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})\leq\beta(1)-\beta(0)<\infty.$
By Proposition C.3, the sequence $(\xi_{n})_{n\in\mathbb{N}}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determind by
$q_{\xi^{\ast}}(u)={\rm TV}_{[0,u]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})\quad\mbox{ for all }u\in(0,1].$
Then for $(u,v)\in\Delta$, we obtain from ${\rm TV}_{[0,v]}={\rm TV}_{[0,u]}+{\rm TV}_{[u,v]}$ that
$q_{\xi^{\ast}}(v)-q_{\xi^{\ast}}^{+}(u)$ $\leq{\rm TV}_{[0,v]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})-{\rm TV}_{[0,u]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})$
$={\rm TV}_{[u,v]}(\{q_{\xi_{n}}\colon n\in\mathbb{N}\})\leq\beta(v)-\beta(u)=\alpha(u,v),$
which shows that $\alpha$ is countable stable. ∎
A set $Q$ of quantile functions is of bounded variation if ${\rm TV}_{[0,1]}(Q)<\infty$.
Lemma 4.14.
Let $A$ be a non-empty subset of $\mathcal{M}_{\infty}$. If $A$ is countably directed w.r.t. $\preccurlyeq_{\rm disp}$, then the set $Q:=\{q_{\xi}\colon\xi\in A\}$ is of bounded variation and satisfies
$\sup_{\xi\in A}(q_{\xi}(v)-q_{\xi}^{+}(u))={\rm TV}_{[u,v]}(Q)\quad\mbox{for all }(u,v)\in\Delta.$
Proof.
By contradiction, suppose that ${\rm TV}_{[0,1]}(Q)=\infty$. Then there exists a sequence $(\xi_{n})_{n\in\mathbb{N}}\in A$ and partitions $\pi_{1}\subset\pi_{2}\subset\cdots$ of $[0,1]$ with $\lim_{n\to\infty}S_{\pi}(q_{\xi_{n}})=\infty$. Since $A$ is countably directed, the sequence $(\xi_{n})_{n\in\mathbb{N}}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$. Then
$\infty=\lim_{n\to\infty}S_{\pi_{n}}(q_{\xi_{n}})\leq\sup_{n\to\infty}S_{\pi_{n}}(q_{\xi^{\ast}})\leq{\rm TV}_{[0,1]}(q_{\xi^{\ast}}),$
in contradiction to ${\rm TV}_{[0,1]}(q_{\xi^{\ast}})<\infty$ since $\xi^{\ast}$ is essentially bounded. This shows that $Q$ is of bounded variation.
Define $\alpha(u,v):=\sup_{\xi\in A}(q_{\xi}(v)-q_{\xi}^{+}(u))$. For $(u,v)\in\Delta$, it follows from Proposition C.2 that
$\alpha(u,v)\leq{\rm TV}_{[u,v]}(Q).$
To prove the equality, suppose by contradiction that $\alpha(u,v)<{\rm TV}_{[u,v]}(Q)$. Then by Proposition C.2, there exists a partition $\pi:=\{u=t_{0}<t_{1}<\cdots<t_{n}=v\}$ with
$\alpha(u,v)<\sum_{i=1}^{n}\sup_{\xi\in A}(q_{\xi}(t_{i})-q_{\xi}^{+}(t_{i-1}))<\sum_{i=1}^{n}(q_{\xi_{i}}(t_{i})-q_{\xi_{i}}^{+}(t_{i-1}))$
for some $\xi_{1},\xi_{2},\cdots,\xi_{n}\in A$. Since $A$ is countably directed, the exists $\xi^{\ast}\in A$ such that $\xi_{1},\dots,\xi_{n}\preccurlyeq_{\rm disp}\xi^{\ast}$. Then
$\alpha(u,v)<\sum_{i=1}^{n}(q_{\xi^{\ast}}(t_{i})-q_{\xi^{\ast}}^{+}(t_{i-1}))\leq q_{\xi^{\ast}}(v)-q_{\xi^{\ast}}^{+}(u)\leq\alpha(u,v),$
which is a contradiction. ∎
For a function $\psi\colon\mathcal{M}_{\infty}\to(-\infty,\infty]$ with sublevel sets $A_{s}:=\{\xi\in\mathcal{M}_{\infty}\colon\psi(\xi)\leq s\}$, we define $\alpha_{\min}(s,u,v):=\sup_{\xi\in A_{s}}(q_{\xi}(v)-q_{\xi}^{+}(u))$ and $\alpha^{+}_{\min}(s,u,v):=\inf_{s<t}\alpha_{\min}(t,u,v)$ for all $s\in\mathbb{R}$ and $(u,v)\in\Delta$.
Let $I:=\{s\in\mathbb{R}\colon A_{s}\neq\emptyset\}$. If $\psi$ has countably directed sublevel sets, it follows from Lemma 4.14 that for every $s\in I$, the set $Q_{s}:=\{q_{\xi}\colon\xi\in A_{s}\}$ is of bounded variation, and $\alpha_{\min}(s,u,v)=\beta(s,v)-\beta(s,u)$ for all $(u,v)\in\Delta$ where $\beta(s,u):={\rm TV}_{[0,u]}(Q_{s})$. Moreover, the mapping $\beta\colon\mathbb{R}\times(0,1)\to\mathbb{R}$ is non-decreasing in the second argument and satisfies $\beta(s,v)-\beta(s,u)\leq\beta(s^{\prime},v)-\beta(s^{\prime},u)$ for all $s\leq s^{\prime}$. As a direct application of Theorem 3.1, we obtain the following representation result for maxitive functions w.r.t. the dispersive stochastic order.
Theorem 4.15.
Let $\psi\colon\mathcal{M}_{\infty}\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$. Then the following conditions are equivalent:
  1. (1)
    $\psi$ is monotone and $(A_{s})_{s\in\mathbb{R}}$ are countably directed upwards w.r.t. $\preccurlyeq_{\rm disp}$.
  2. (2)
    There exists a function $\beta\colon I\times(0,1)\to\mathbb{R}$ which is non-decreasing in the second argument, and satisfies $\beta(s,v)-\beta(s,u)\leq\beta(s^{\prime},v)-\beta(s^{\prime},u)$ for all $s\leq s^{\prime}$ and $(u,v)\in\Delta$. Moreover, it holds
    $\psi(\xi)=\inf\left\{s\in\mathbb{R}\colon\sup_{(u,v)\in\Delta}\{q_{\xi}(v)-q_{\xi}(u^{+})-\alpha(s,u,v)\}\leq 0\right\}\quad\mbox{ for all }\xi\in\mathcal{M}_{\infty},$
    where $\alpha(s,u,v):=\beta(s,v)-\beta(s,u)$ for all $s\in I$, and $\alpha(s,u,v):=\infty$ otherwise. The infimum is attained for all $\xi\in\mathcal{M}_{\infty}$ with $\psi(\xi)<\infty$.
  3. (3)
    For every $s\in I$, the set $Q_{s}:=\{q_{\xi}\colon\xi\in A_{s}\}$ is of bounded variation and $\alpha_{\min}(s,u,v)={\rm TV}_{[u,v]}(Q_{s})$ for all $(u,v)\in\Delta$. Moreover, it holds
    $\psi(\xi)=\sup_{(u,v)\in\Delta}G(q_{\xi}(v)-q_{\xi}^{+}(u),u,v)\quad\mbox{ forall }\xi\in\mathcal{M}_{\infty},$
    where $G(t,u,v):=\inf\{s\in\mathbb{R}\colon t\leq\alpha_{\min}^{+}(s,u,v)\}$ for all $(t,u,v)\in\mathbb{R}\times\Delta$.
Appendix A Auxiliary results for maxitive risk measures on $C_{b}(S)$
Let $(S,d)$ be a Polish space, and denote by $\mathcal{X}$ a convex cone of functions $f\colon S\to\mathbb{R}$ with $C_{b}(S)\subset\mathcal{X}$. Here, $C_{b}(S)$ denotes the space of all bounded continuous functions. The set $\mathcal{X}$ is endowed with the pointwise order. For a function $\psi\colon\mathcal{X}\to(-\infty,\infty]$, we frequently deal with the following properties:
  • (T)
    $\psi(f+c)=\psi(f)+c$ for all $c\in\mathbb{R}$.
  • (M)
    $\psi(f)\leq\psi(g)$ whenever $f\leq g$.
  • (C)
    $\psi(\lambda f+(1-\lambda)g)\leq\lambda\psi(f)+(1-\lambda)\psi(g)$ for all $\lambda\in[0,1]$.
The function $\psi$ is continuous from below if $\psi(f)=\lim_{n\to\infty}\psi(f_{n})$ for all sequences $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{X}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ and $f=\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{X}$. Moreover, the function $\psi$ is countably maxitive if $\psi(f)=\sup_{n\in\mathbb{N}}\psi(f_{n})$ for all sequences $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{X}$ with $f=\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{X}$.
For $\psi\colon\mathcal{X}\to(-\infty,\infty]$, we consider the acceptance set $\mathcal{A}:=\{f\in\mathcal{X}\colon\psi(f)\leq 0\}$, and define the convex conjugate $\psi_{C_{b}(S)}^{\ast}\colon{\rm ca}_{1}^{+}(S)\to[0,\infty]$ by
$\psi^{\ast}_{C_{b}(S)}(\mu):=\sup_{f\in C_{b}(S)}\left\{\int f\,{\rm d}\mu-\psi(f)\right\}.$
Here, ${\rm ca}_{1}^{+}(S)$ denotes the set of all Borel probability measures on $S$. In the particular case of a Dirac measure $\delta_{x}$, we get
$I_{\min}(x):=\psi^{\ast}_{C_{b}(S)}(\delta_{x})=\sup_{f\in C_{b}(S)}\{f(x)-\psi(f)\quad\mbox{for all }x\in S.$
The following result was recently shown by Delbaen [4, 5], see also [15] for an alternative proof.
Theorem A.1.
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T), (M) and (C). Then the following statements are equivalent:
  • (i)
    $\psi$ is continuous from below.
  • (ii)
    $\psi$ admits the representation
    $\psi(f)=\sup_{\mu\in{\rm ca}_{1}^{+}(S)}\left\{\int f\,{\rm d}\mu-\psi_{C_{b}(S)}^{\ast}(\mu)\right\}.$
In combination with [12, Theorem 2.1], we obtain the following representation result.
Theorem A.2.
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T) and (M). Then the following statements are equivalent:
  • (i)
    $\psi$ is countably maxitive.
  • (ii)
    $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$ for every sequence $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$.
  • (iii)
    $\psi$ admits the representation
    $\psi(f)=\sup_{x\in S}\left\{f(x)-I_{\min}(x)\right\}\quad\mbox{for all }f\in C_{b}(S).$
  • (iv)
    $\psi$ satisfies (C) and is continuous from below. Moreover,
    $\psi_{C_{b}(S)}^{\ast}(\mu)=\int I_{\min}(x)\,\mu({\rm d}x)\quad\mbox{for all }\mu\in{\rm ca}_{1}^{+}(S).$
  • (v)
    There exists a function $\gamma\colon S\to[0,\infty]$ such that
    $\mathcal{A}=\{f\in C_{b}(S)\colon f\leq\gamma\}.$
Proof.
The argumentation is analogous to that in [12, Theorem 2.1]. Similar to the proof of [12, Theorem 2.1], it holds
[…]
for all $\mu\in{\rm ca}_{1}^{+}(S)$, where ${\rm ess.sup}_{\mu}\mathcal{A}$ denotes the essential supremum of $\mathcal{A}$ w.r.t. $\mu$.
$(i)\Rightarrow(ii)$: Let $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since $\psi$ is countably maxitive, it holds $\psi(\sup_{n\in\mathbb{N}}f_{n})=\sup_{n\in\mathbb{N}}\psi(f_{n})\leq 0$, which shows that $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$.
$(ii)\Rightarrow(i)$: Let $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since by (T), $f_{n}-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$ for all $n\in\mathbb{N}$, we obtain $f-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$. By (T) and (M), it holds
$\psi(f)\leq\sup_{n\in\mathbb{N}}\psi(f_{n})\leq\psi(f).$
$(i)\Rightarrow(iv):$ Since $\psi$ is countably maxitive, by [12, Proposition 2.1], it satisfies (C). Fix $\mu\in{\rm ca}_{1}^{+}(S)$. Since $\mathcal{A}$ is directed upwards, there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ such that $f_{n}\nearrow{\rm ess.sup}_{\mu}\mathcal{A}$ $\mu$-almost surely. By (A.1), we have that ${\rm ess.sup}_{\mu}\mathcal{A}=I_{\min}$ $\mu$-almost surely. Hence, by monotone convergence,
$\psi_{C_{b}(S)}^{\ast}(\mu)=\sup_{f\in\mathcal{A}}\int f\,{\rm d}\mu\geq\sup_{n\in\mathbb{N}}\int f_{n}\,{\rm d}\mu=\int I_{\min}(x)\,\mu({\rm d}x).$
Moreover,
$\psi_{C_{b}(S)}^{\ast}(\mu)=\sup_{f\in\mathcal{A}}\int f\,{\rm d}\mu\leq\int{\rm ess.sup}_{\mu}\mathcal{A}\,{\rm d}\mu=\int I_{\min}(x)\,\mu({\rm d}x).$
$(iv)\Rightarrow(iii):$ For every $f\in C_{b}(S)$, it follows from Theorem A.1 that
$\psi(f)$ $=\sup_{\mu\in{\rm ca}_{1}^{+}(S)}\left\{\int f\,{\rm d}\mu-\psi_{C_{b}(S)}^{\ast}(\mu)\right\}=\sup_{\mu\in{\rm ca}_{1}^{+}(S)}\left\{\int f\,{\rm d}\mu-\int I_{\min}(x)\,\mu({\rm d}x)\right\}$
$=\sup_{\mu\in{\rm ca}_{1}^{+}(S)}\left\{\int f(x)-I_{\min}(x)\,\mu({\rm d}x)\right\}=\sup_{x\in S}\left\{f(x)-I_{\min}(x)\right\}.$
$(iii)\Rightarrow(v):$ From the representation, $f\in\mathcal{A}$ if and only if $f(x)\leq I_{\min}(x)$ for all $x\in S$. Hence, $\mathcal{A}=\{f\in C_{b}(S)\colon f\leq\gamma\}$ with $\gamma=I_{\min}$.
$(v)\Rightarrow(ii):$ Let $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since $f_{n}\leq\gamma$ for all $n\in\mathbb{N}$, it follows that $\sup_{n\in\mathbb{N}}f_{n}\leq\gamma$. This shows that $f\in\mathcal{A}$. ∎
We denote by $L(S)$ the set of all lower semicontinuous functions $f\colon S\to\mathbb{R}$.
Corollary A.3.
Suppose that $\psi\colon{L}(S)\to(-\infty,\infty]$ satisfies $\psi(0)\in\mathbb{R}$, (T), (M) and is countably maxitive. Then
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\mbox{for all }f\in{L}(S).$
Proof.
Since $\psi(0)\in\mathbb{R}$, it follows from (T) that the restriction of $\psi$ to $C_{b}(S)$ is real-valued. Hence, it follows from Theorem A.2 that
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\mbox{for all }f\in C_{b}(S).$
Now, let $f\in L(S)$ being bounded from below. Then there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f_{n}\nearrow f$. For instance, $f_{n}(x):=\inf\left\{f(y)+nd(x,y)\colon y\in S\right\}\wedge n$ is as desired. Then since $\psi$ is continuous from below,
$\psi(f)$ $=\sup_{n\in\mathbb{N}}\psi(f_{n})=\sup_{n\in\mathbb{N}}\sup_{x\in S}\{f_{n}(x)-I_{\min}(x)\}$
$=\sup_{x\in S}\sup_{n\in\mathbb{N}}\{f_{n}(x)-I_{\min}(x)\}=\sup_{x\in S}\{f(x)-I_{\min}(x)\}.$
Finally, for general $f\in L(S)$, since $\psi$ is countably maxitive, for every $n\in\mathbb{N}$, we obtain
$\psi(f)\vee\psi(-n)$ $=\psi(f\vee(-n))$
$=\sup_{x\in S}\{f(x)\vee(-n)-I_{\min}(x)\}$
$=\sup_{x\in S}\{(f(x)-I_{\min}(x))\vee(-n-I_{\min}(x))\}.$
Letting $n\to\infty$, we obtain $\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}$. ∎
Appendix B Auxiliary results for quantile functions
Suppose that $(\Omega,\mathcal{F},\mathbb{P})$ is an atomless probability space. Recall that $\mathcal{M}$ denotes the set of all random variables $\xi\colon\Omega\to\mathbb{R}$, where two of them are identified if they have the same law. Given $\xi\in\mathcal{M}$, we define the cumulative distribution function by $F_{\xi}(x):=\mathbb{P}(\xi\leq x)$ and set $F_{\xi}(x^{-})=\sup_{y<x}F_{\xi}(y)$ for all $x\in\mathbb{R}$. Moreover, the quantile function $q_{\xi}\colon(0,1)\to\mathbb{R}$ and upper quantile function $q_{\xi}^{+}\colon(0,1)\to\mathbb{R}$ are defined by
$q_{\xi}(t):=\inf\{x\in\mathbb{R}\colon t\leq F_{\xi}(x)\}\quad\mbox{and}\quad q_{\xi}^{+}(t):=\inf\{x\in\mathbb{R}\colon t<F_{\xi}(x)\}.$
Lemma B.1.
For all $\xi\in\mathcal{M}$ and $t\in(0,1)$, the following statements hold:
  • (i)
    $[q_{\xi}(t),\infty)=\{x\in\mathbb{R}\colon t\leq F_{\xi}(x)\}$.
  • (ii)
    $(-\infty,q^{+}_{\xi}(t)]=\{x\in\mathbb{R}\colon t\geq F_{\xi}(x^{-})\}$.
  • (iii)
    $\sup_{s<t}q_{\xi}(s)=\sup_{s<t}q^{+}_{\xi}(s)=q_{\xi}(t)$.
  • (iv)
    $\inf_{s>t}q_{\xi}(s)=\inf_{s>t}q^{+}_{\xi}(s)=q^{+}_{\xi}(t)$.
  • (v)
    The function $q_{\xi}\colon(0,1)\to\mathbb{R}$ is left-continuous.
  • (vi)
    The function $q_{\xi}^{+}\colon(0,1)\to\mathbb{R}$ is right-continuous.
  • (vii)
    The function $(s,r)\mapsto q_{\xi}(r)-q_{\xi}^{+}(s)$ is lower semicontinuous on $(0,1)^{2}$.
Proof.
(i) and (ii) follow from the definitions of $q$ and $q^{+}$, using that $F_{\xi}$ is non-decreasing and right-continuous.
For $a:=\sup_{s<t}q_{\xi}(s)$, it follows from (i) that
$[a,\infty)=\underset{s<t}{\bigcap}[q_{\xi}(s),\infty)=\underset{s<t}{\bigcap}F_{\xi}^{-1}([s,\infty)])=F_{\xi}^{-1}([t,\infty)])=[q_{\xi}(t),\infty),$
where $F_{\xi}^{-1}([s,\infty))=\{x\in\mathbb{R}\colon s\leq F_{\xi}(x)\}$. This shows $\sup_{s<t}q_{\xi}(s)=q_{\xi}(t)$. For $s<t$, it holds $q_{\xi}(s)\leq q_{\xi}^{+}(s)\leq q_{\xi}(t)$, so that $q_{\xi}(t)=\sup_{s<t}q_{\xi}(s)\leq\sup_{s<t}q_{\xi}^{+}(s)\leq q_{\xi}(t)$, which shows (iii).
Define $G(s):=F_{\xi}(s^{-})$ for all $s\in\mathbb{R}$. For $b:=\inf_{s>t}q_{\xi}^{+}(s)$, it follows from (ii) that
$(-\infty,b]=\underset{s>t}{\bigcap}(-\infty,q_{\xi}^{+}(s)]=\underset{s>t}{\bigcap}G^{-1}((-\infty,s])=G^{-1}((-\infty,t])=(-\infty,q_{\xi}^{+}(t)],$
where $G^{-1}([s,\infty))=\{x\in\mathbb{R}\colon s\leq G(x)\}$. This shows $\inf_{s>t}q_{\xi}^{+}(s)=q_{\xi}^{+}(t)$. For $s>t$, it holds $q_{\xi}^{+}(s)\geq q_{\xi}(s)\geq q_{\xi}^{+}(t)$, so that $q_{\xi}^{+}(t)=\inf_{s>t}q_{\xi}^{+}(s)\geq\inf_{s>t}q_{\xi}(s)\geq q_{\xi}^{+}(t)$, which shows (iv).
Finally, (v), (vi) and (vii) are consequences of (iii) and (iv). ∎
Lemma B.2.
For $\xi,\eta\in\mathcal{M}$, the following statements are equivalent:
  • (i)
    $q_{\xi}(u)-q_{\xi}(t)\leq q_{\eta}(u)-q_{\eta}(t)$ for all $10<t\leq u<$ 1.
  • (ii)
    $q_{\xi}(u)-q_{\xi}^{+}(t)\leq q_{\eta}(u)-q_{\eta}^{+}(t)$ for all $10<t<u<$ 1.
Proof.
$(ii)\Rightarrow(i)$: Let $10<t\leq u<$ 1. For $s\in(0,t)$, it holds $q_{\xi}(u)-q_{\xi}^{+}(s)\leq q_{\eta}(u)-q_{\eta}^{+}(s)$. Letting $s\to t^{-}$, it follows from (iii) in Lemma B.1 that
$q_{\xi}(u)-q_{\xi}(t)\leq q_{\eta}(u)-q_{\eta}(t).$
$(i)\Rightarrow(ii)$: Let $10<t<u<$ 1. For $s\in(t,u)$, it holds $q_{\xi}(u)-q_{\xi}(s)\leq q_{\eta}(u)-q_{\eta}(s)$. Letting $s\to t^{+}$, it follows from (iv) in Lemma B.1 that $q_{\xi}(u)-q_{\xi}^{+}(t)\leq q_{\eta}(u)-q_{\eta}^{+}(t)$. ∎
Lemma B.3.
If $q\colon(0,1)\to\mathbb{R}$ is non-decreasing and left-continuous, then there exists $\xi\in\mathcal{M}$ such that $q(u)=q_{\xi}(u)$ for all $u\in(0,1)$.
Proof.
If $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, then there exists a uniformly distributed random variable $U$ with values in $(0,1)$ defined on $(\Omega,\mathcal{F},\mathbb{P})$, see e.g. [3, Theorem 1] and [9, Proposition A.27]. Then $\xi=q(U)$ is as desired. To that end, define $F\colon\mathbb{R}\to[0,1]$ by
$F(x):=0\vee\sup\{s\in(0,1)\colon x\geq q(s)\},$
which satisfies
  • (1)
    $q(F(x))\leq x$ for all $x\in\mathbb{R}$ (where we set $q(0):=q(0^{+})$ and $q(1):=q(1^{-})$),
  • (2)
    $F(q(s))\geq s$ for all $s\in(0,1)$.
For $x\in\mathbb{R}$, it follows from (2) that $\{q(U)\leq x\}\subset\{F(q(U))\leq F(x)\}\subset\{U\leq F(x)\}$. Hence, for every $u\in(0,1)$,
$q_{\xi}(u)$ $=\inf\{x\in\mathbb{R}\colon\mathbb{P}(q(U)\leq x)\geq u\}$
$\geq\inf\{x\in\mathbb{R}\colon\mathbb{P}(U\leq F(x))\geq u\}$
$=\inf\{x\in\mathbb{R}\colon F(x)\geq u\}.$
Let $a:=\inf\{x\in\mathbb{R}\colon F(x)\geq u\}$, and $x_{n}\downarrow a$ with $F(x_{n})\geq u$. It follows from (1) that $x_{n}\geq q(u)$, and therefore $q_{\xi}(u)\geq a\geq q(u)$ by letting $n\to\infty$.
As for the other inequality, we have that $\mathbb{P}(q(U)\leq q(u))\geq\mathbb{P}(U\leq u)=u$. Hence, it follows from the definition of $q_{\xi}(u)$ that $q_{\xi}(u)\leq q(u)$. ∎
Recall that for $\xi\in\mathcal{M}_{1}$, the integrated quantile function $Q_{\xi}\colon(0,1)\to\mathbb{R}$ is defined by
$Q_{\xi}(u):=\int_{u}^{1}q_{\xi}(v)\,{\rm d}v.$
Lemma B.4.
If $Q\colon(0,1)\to\mathbb{R}$ is concave and satisfies $Q(0^{+})\in\mathbb{R}$ and $Q(1^{-})=0$, then there exists $\xi\in\mathcal{M}_{1}$ such that $Q(u)=Q_{\xi}(u)$ for all $u\in(0,1)$.
Proof.
Since $Q$ is concave, there exist a non-decreasing function $q\colon(0,1)\to\mathbb{R}$ and $c\in(0,1)$ such that
$Q(u)=Q(c)-\int_{c}^{u}q(v){\rm d}v,$
see e.g. [21, Theorem 12A]. By replacing $q$ by its left-continuous modification, we can assume that $q$ is left-continuous. Using that $Q(1^{-})=0$, we get that
$Q(u)=\int_{u}^{1}q(v)\,{\rm d}v.$
Since $q$ is left-continuous, it follows from Lemma B.3 that $q=q_{\xi}$ for some $\xi\in\mathcal{M}$. It remains to show that $\xi\in\mathcal{M}_{1}$. We have
$\mathbb{E}_{\mathbb{P}}[|\xi|]$ $=\int_{0}^{1}|q_{\xi}(v)|{\rm d}v$
$=-\int_{0}^{F_{\xi}(0)}q_{\xi}(v){\rm d}v+\int_{F_{\xi}(0)}^{1}q_{\xi}(v){\rm d}v$
$=-Q(0^{+})+2Q(F_{\xi}(0))<\infty.$
Hence, $\xi$ has finite first moment, and consequently $\xi\in\mathcal{M}_{1}$. ∎
Appendix C Auxiliary results for the total variation
Suppose that $(\Omega,\mathcal{F},\mathbb{P})$ is an atomless probability space. Recall that for a non-empty set $Q$ of quantile functions and $[u,v]\subset[0,1]$, we define its total variation on $[u,v]$ as
${\rm TV}_{[u,v]}\left(Q\right):=\sup_{\pi\in\Pi[u,v]}S_{\pi}(Q),$
where $\Pi[u,v]$ consists of all partitions $\pi=\{u=t_{0}<t_{1}<\cdots<t_{n}=v\}$ for some $n\in\mathbb{N}$. Here, we set $S_{\pi}(Q):=\sum_{i=1}^{n}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))$ with $q(0):=q(0^{+})$ and $q(1):=q(1^{-})$.
We say that the set $Q$ is of bounded variation if ${\rm TV}_{[0,1]}(Q)<\infty$. It is straightforward to verify that ${\rm TV}_{[0,v]}={\rm TV}_{[0,u]}+{\rm TV}_{[u,v]}$ for all $10<u\leq v<$ 1.
Proposition C.1.
The mapping $t\mapsto{\rm TV}_{[0,t]}(Q)$ is non-decreasing and left-continuous on $(0,1]$.
Proof.
By definition, the mapping $t\mapsto{\rm TV}_{[0,t]}(Q)$ is non-decreasing. Given $t\in(0,1]$, we fix a partition $\pi=\{t_{0}<t_{1}<\cdots<t_{n}\}$ of $[0,t]$. For every $u\in(0,t)$, we consider the partition $\pi_{u}:=(\pi\cap[0,u])\cup\{u\}\in\Pi[0,u]$. Then since every $q\in Q$ is non-decreasing and left-continuous, we have that
$\sup_{u<t}S_{\pi_{u}}(Q)$ $=\sup_{u\in(t_{n-1},t)}S_{\pi_{u}}(Q)$
$=\sum_{i=1}^{n-1}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))+\sup_{u\in(t_{n-1},t)}\sup_{q\in Q}(q(u)-q(t_{n-1}))$
$=\sum_{i=1}^{n-1}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))+\sup_{q\in Q}\sup_{u\in(t_{n-1},t)}(q(u)-q(t_{n-1}))$
$=\sum_{i=1}^{n-1}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))+\sup_{q\in Q}(q(t)-q(t_{n-1}))$
$=S_{\pi}(Q).$
Since ${\rm TV}_{[0,u]}(Q)=\sup_{\pi\in\Pi[0,u]}S_{\pi}(Q)=\sup_{\pi\in\Pi[0,t]}S_{\pi_{u}}(Q)$ for all $u<t$, it follows that
$\lim_{u\uparrow t}{\rm TV}_{[0,u]}(Q)$ $=\sup_{u<t}\sup_{\pi\in\Pi[0,t]}S_{\pi_{u}}(Q)=\sup_{\pi\in\Pi[0,t]}\sup_{u<t}S_{\pi_{u}}(Q)$
$=\sup_{\pi\in\Pi[0,t]}S_{\pi}(Q)={\rm TV}_{[0,t]}(Q).\qed$
For $[u,v]\subset[0,1]$ and $\pi\in\Pi[u,v]$, we define $\underline{S}_{\pi}(Q):=\sum_{i=1}^{n}\sup_{q\in Q}(q_{\xi}(t_{i})-q_{\xi}^{+}(t_{i-1}))$.
Proposition C.2.
For every $[u,v]\subset[0,1]$, it holds
${\rm TV}_{[u,v]}(Q)=\sup_{\pi\in\Pi[u,v]}\underline{S}_{\pi}(Q).$
Proof.
By contradiction, we assume that $\sup_{\pi\in\Pi[u,v]}S_{\pi}(Q)>\sup_{\pi\in\prod[u,v]}\underline{S}_{\pi}(Q)$. Then, there exists $\bar{\pi}=\{u=t_{0}<t_{1}<\cdots<t_{n}=v\}\in\Pi[u,v]$ with
$S_{\bar{\pi}}(Q)>\sup_{\pi\in\Pi[0,t]}\underline{S}_{\pi}(Q).$
Fix $\varepsilon>0$ such that $S_{\bar{\pi}}(Q)-\varepsilon>\sup_{\pi\in\Pi[0,t]}\underline{S}_{\pi}(Q)$. For every $i\in\{1,2,\cdots,n\}$, the mapping $t\mapsto\sup_{q\in Q}(q(t)-q(t_{i-1}))$ is left-continuous. Hence, there exists $u_{i}\in[t_{i-1},t_{i})$ such that
$\sup_{q\in Q}(q(u_{i})-q(t_{i-1}))>\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))-\frac{\varepsilon}{n}.$
Since $q^{+}(u_{i})\leq q(t_{i})$ for all $i$, we get
$\sup_{q\in Q}(q(u_{i})-q^{+}(u_{i-1}))\geq\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))-\frac{\varepsilon}{n}\quad\mbox{for all }i\in\{1,2,\cdots,n\},$
where $u_{0}:=t_{0}=u$. For the partition $\pi^{\prime}:=\{u=u_{0}<\cdots<u_{n}<t_{n}=v\}\in\Pi[u,v]$, we obtain
$\underline{S}_{\pi^{\prime}}(Q)\geq S_{\bar{\pi}}(Q)-\varepsilon>\sup_{\pi\in\Pi[0,t]}\underline{S}_{\pi}(Q),$
which is a contradiction. ∎
The following two results are formulated in the setting of Subsection 4.5.
Proposition C.3.
Suppose that $Q:=\{q_{\xi}\colon\xi\in A\}$ is of bounded variation for some $A\subset\mathcal{M}$. Then there exists $\xi^{\ast}\in\mathcal{M}$ such that $q_{\xi^{\ast}}(t)={\rm TV}_{[0,t]}(Q)$ for all $t\in(0,1]$. Moreover, $\xi^{\ast}$ is a supremum of $A$ w.r.t. $\preccurlyeq_{\rm disp}$.
Proof.
By Proposition C.1 and the fact that $Q$ is of bounded variation, the function $t\mapsto{\rm TV}_{[0,t]}(Q)$ is real-valued, non-decreasing and left-continuous. Hence, there exists $\xi^{\ast}\in\mathcal{M}$ such that $q_{\xi^{\ast}}(t)={\rm TV}_{[0,t]}(Q)$ for all $t\in(0,1]$. Given $u,v\in\mathbb{R}$ with $10<u\leq v<$ 1, for every $\xi\in A$, we have that
$q_{\xi}(v)-q_{\xi}(u)\leq{\rm TV}_{[u,v]}(Q)={\rm TV}_{[0,v]}(Q)-{\rm TV}_{[0,u]}(Q)=q_{\xi^{\ast}}(v)-q_{\xi^{\ast}}(u).$
Thus, $\xi^{\ast}$ is an upper bound of $A$ w.r.t. $\preccurlyeq_{\rm disp}$. On the other hand, if $\eta$ is another upper bound of $A$ w.r.t. $\preccurlyeq_{\rm disp}$, we necessarily have that
$q_{\xi^{\ast}}(v)-q_{\xi^{\ast}}(u)={\rm TV}_{[u,v]}(Q)\leq{\rm TV}_{[u,v]}(q_{\eta})=q_{\eta}(v)-q_{\eta}(u),$
where we have used in the last equality that $q_{\eta}$ is non-decreasing. This implies that $\xi^{\ast}\preccurlyeq_{\rm disp}\eta$. ∎
Corollary C.4.
The partially ordered space $(\mathcal{M}_{\infty}/_{\sim},\preccurlyeq_{\rm disp})$ is a complete lattice.
Proof.
Let $A\subset\mathcal{M}_{\infty}/_{\sim}$ be a set which is bounded from above w.r.t. $\preccurlyeq_{\rm disp}$, and denote by $A_{0}$ the set of respective representatives $\xi$ with $q_{\xi}(0^{+})=0$. Suppose that $\xi^{\ast}$ is an upper bound of $A_{0}$. Then
${\rm TV}_{[0,1]}(A_{0})\leq{\rm TV}_{[0,1]}(q_{\xi^{\ast}})=q_{\xi^{\ast}}(1^{-})-q_{\xi^{\ast}}(0^{+})<\infty.$
Hence, $A_{0}$ is of bounded variation and has a supremum in $\mathcal{M}_{\infty}$ w.r.t. $\preccurlyeq_{\rm disp}$ by Proposition C.3. Then passing to the quotient, $A$ has a unique supremum in $\mathcal{M}_{\infty}/_{\sim}$. ∎